2.7. Suppose that we are testing Ho : μ1 Ha versus Ho: > μ2 where the two sample sizes are ni n, 10. Both sample variances are unknown but assumed equal. Find bounds on the P-value for the following observed values of the test statistic. (a) to= 2.31 (b) toー3.60 (c) to-1.95 (d) 0-219 2.8. Consider the following sample data: 9.37, 13.04, 11.69, 8.21, 11.18, 10.41, 13.15, 11.51, 13.21, and 7.75. Is it reasonable to assume that this data is a sample from a normal distribution? Is there evidence to support a claim that the mean of the population is 10? 2.9. put for a hypothesis-testing problem A computer program has produced the following out- Difference in sample means: 2.35 Degrees of freedom: 18 Standard error of the difference in sample means: ? Test statistic: to2.01 P-value: 0.0298 (a) What is the missing value for the standard error? (b) Is this a two-sided or a one-sided test? (c) If α 0.05, what are your conclusions? (d) Find a 90% two-sided CI on the difference in means

Please solve 2.7, 2.8 and 2.9 problems.
Answer #1

Solution:

    27) We are given that:

H_{0} : \mu_{1} = \mu_{2} Vs H_{1} : \mu_{1} > \mu_{2}

Sample sizes are n1 = n2 = 10

Variances are unknown but assumed equal.
We have to find bounds on P-value for the following observed values of the test statistic.

Since H_{1} : \mu_{1} > \mu_{2} is right tailed , we use one tail area for finding P-value intervals.

df = n1 + n2 - 2 = 10 + 10 - 2 = 18

So we look in t table for row of df = 18 and find the interval in which t test statistic fall.

Part a) t0 = 2.31

Look in t table for df row = 18 and find the interval in which t = 2.31 fall , then find corresponding one tail area interval, which would be bounds on P-value.

t0 = 2.31 fall in between t = 2.101 to t = 2.552, thus corresponding one tail area is in between 0.01 to 0.025

Thus bounds on P-value are:

0.01 < P-value < 0.025

Part b) t0 = 3.60

Look in t table for df row = 18 and find the interval in which t = 3.60 fall , then find corresponding one tail area interval, which would be bounds on P-value.

t Table cum. prob one-tail0.50 0.25 0.20 0.15 0.10 0.05 0.025 0.00.005 .001 0.0005 two-tails1.00 0.50 0.40 0.30 0.20 0.10 0.0

t0 = 3.60 fall in between t = 2.878 to t = 3.610, thus corresponding one tail area is in between 0.001 to 0.005

Thus bounds on P-value are:

0.001 < P-value < 0.005

Part c) t0 = 1.95

Look in t table for df row = 18 and find the interval in which t = 1.95 fall , then find corresponding one tail area interval, which would be bounds on P-value.

t Table cum. prob 2 one-tail0.50 0.25 0.20 0.15 0.10 0.05 0.025 0.0 0.005 0.0010.0005 2 1 0 2 0 1 0.000 000 1.376 1.963 3.078

t0 = 1.95 fall in between t = 1.734 to t = 2.101, thus corresponding one tail area is in between 0.025 to 0.05

Thus bounds on P-value are:

0.025 < P-value < 0.05

Part d) t0 = 2.19

Look in t table for df row = 18 and find the interval in which t = 2.19 fall , then find corresponding one tail area interval, which would be bounds on P-value.

t0 = 2.19 fall in between t = 2.101 to t = 2.552, thus corresponding one tail area is in between 0.01 to 0.025

Thus bounds on P-value are:

0.01 < P-value < 0.025